Solve.
If the hypotenuse of a right triangle is 12 feet and a side is 6 feet, what is the length of the other side. Express your answer in radical form.

Answer by checkley71(8403)

6^2+X^2=12^2
36+X^2=144
X^2=144-36
X^2=108
X=SQRT(108)